1. what is the volume when one mole of different liquids and solids? and why?
1
Expert's answer
2013-01-09T08:21:32-0500
v - Amount of substance
m - mass
M - Molar mass
v=m/M
m=v*M
V=m/density
V - volume
V=v*M/density
if v=1
V=M/density
the volume when one mole of differentliquids and solids depends of their molar masses and densities